\begin{thebibliography}{26}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Acharya et~al.(2015)Acharya, Sudarshan, Adeli, Santhosh, Koh,
  Puthankatti, and Adeli]{acharya2015novel}
U~Rajendra Acharya, Vidya~K Sudarshan, Hojjat Adeli, Jayasree Santhosh, Joel~EW
  Koh, Subha~D Puthankatti, and Amir Adeli.
\newblock A novel depression diagnosis index using nonlinear features in eeg
  signals.
\newblock \emph{European neurology}, 74\penalty0 (1-2):\penalty0 79--83, 2015.

\bibitem[Acharya et~al.(2018)Acharya, Oh, Hagiwara, Tan, Adeli, and
  Subha]{acharya2018automated}
U~Rajendra Acharya, Shu~Lih Oh, Yuki Hagiwara, Jen~Hong Tan, Hojjat Adeli, and
  D~Puthankattil Subha.
\newblock Automated eeg-based screening of depression using deep convolutional
  neural network.
\newblock \emph{Computer methods and programs in biomedicine}, 161:\penalty0
  103--113, 2018.

\bibitem[Ahmadlou et~al.(2012)Ahmadlou, Adeli, and
  Adeli]{ahmadlou2012fractality}
Mehran Ahmadlou, Hojjat Adeli, and Amir Adeli.
\newblock Fractality analysis of frontal brain in major depressive disorder.
\newblock \emph{International Journal of Psychophysiology}, 85\penalty0
  (2):\penalty0 206--211, 2012.

\bibitem[Ahmadlou et~al.(2013)Ahmadlou, Adeli, and
  Adeli]{ahmadlou2013spatiotemporal}
Mehran Ahmadlou, Hojjat Adeli, and Amir Adeli.
\newblock Spatiotemporal analysis of relative convergence of eegs reveals
  differences between brain dynamics of depressive women and men.
\newblock \emph{Clinical EEG and neuroscience}, 44\penalty0 (3):\penalty0
  175--181, 2013.

\bibitem[Bairy et~al.(2017)Bairy, Lih, Hagiwara, Puthankattil, Faust, Niranjan,
  and Acharya]{bairy2017automated}
G~Muralidhar Bairy, Oh~Shu Lih, Yuki Hagiwara, Subha~D Puthankattil, Oliver
  Faust, UC~Niranjan, and U~Rajendra Acharya.
\newblock Automated diagnosis of depression electroencephalograph signals using
  linear prediction coding and higher order spectra features.
\newblock \emph{Journal of Medical Imaging and Health Informatics}, 7\penalty0
  (8):\penalty0 1857--1862, 2017.

\bibitem[Cai et~al.(2018)Cai, Han, Chen, Sha, Wang, Hu, Yang, Feng, Ding, Chen,
  et~al.]{cai2018pervasive}
Hanshu Cai, Jiashuo Han, Yunfei Chen, Xiaocong Sha, Ziyang Wang, Bin Hu, Jing
  Yang, Lei Feng, Zhijie Ding, Yiqiang Chen, et~al.
\newblock A pervasive approach to eeg-based depression detection.
\newblock \emph{Complexity}, 2018, 2018.

\bibitem[Cukic et~al.(2018)Cukic, Pokrajac, Stokic, Radivojevic,
  Ljubisavljevic, et~al.]{cukic2018eeg}
Milena Cukic, David Pokrajac, Miodrag Stokic, Vlada Radivojevic, Milos
  Ljubisavljevic, et~al.
\newblock Eeg machine learning with higuchi fractal dimension and sample
  entropy as features for successful detection of depression.
\newblock \emph{arXiv preprint arXiv:1803.05985}, 2018.

\bibitem[Erguzel et~al.(2015)Erguzel, Tas, and Cebi]{erguzel2015wrapper}
Turker~Tekin Erguzel, Cumhur Tas, and Merve Cebi.
\newblock A wrapper-based approach for feature selection and classification of
  major depressive disorder--bipolar disorders.
\newblock \emph{Computers in biology and medicine}, 64:\penalty0 127--137,
  2015.

\bibitem[Erguzel et~al.(2016)Erguzel, Sayar, and Tarhan]{erguzel2016artificial}
Turker~Tekin Erguzel, Gokben~Hizli Sayar, and Nevzat Tarhan.
\newblock Artificial intelligence approach to classify unipolar and bipolar
  depressive disorders.
\newblock \emph{Neural Computing and Applications}, 27\penalty0 (6):\penalty0
  1607--1616, 2016.

\bibitem[Faust et~al.(2014)Faust, Ang, Puthankattil, and
  Joseph]{faust2014depression}
Oliver Faust, Peng Chuan~Alvin Ang, Subha~D Puthankattil, and Paul~K Joseph.
\newblock Depression diagnosis support system based on eeg signal entropies.
\newblock \emph{Journal of mechanics in medicine and biology}, 14\penalty0
  (03):\penalty0 1450035, 2014.

\bibitem[Gao et~al.(2018)Gao, Calhoun, and Sui]{gao2018machine}
Shuang Gao, Vince~D Calhoun, and Jing Sui.
\newblock Machine learning in major depression: From classification to
  treatment outcome prediction.
\newblock \emph{CNS neuroscience \& therapeutics}, 24\penalty0 (11):\penalty0
  1037--1052, 2018.

\bibitem[Hossain et~al.(2014)Hossain, Kabir, and Shahjahan]{hossain2014feature}
Md~Zakir Hossain, Md~Monirul Kabir, and Md~Shahjahan.
\newblock Feature selection of eeg data with neuro-statistical method.
\newblock In \emph{2013 International Conference on Electrical Information and
  Communication Technology (EICT)}, pages 1--6. IEEE, 2014.

\bibitem[Hosseinifard et~al.(2013)Hosseinifard, Moradi, and
  Rostami]{hosseinifard2013classifying}
Behshad Hosseinifard, Mohammad~Hassan Moradi, and Reza Rostami.
\newblock Classifying depression patients and normal subjects using machine
  learning techniques and nonlinear features from eeg signal.
\newblock \emph{Computer methods and programs in biomedicine}, 109\penalty0
  (3):\penalty0 339--345, 2013.

\bibitem[Hui et~al.(2018)Hui, Zhang, Chuchen, Jinling, Meijuan, and
  Yiru]{hui2018analysis}
SHEN Hui, Li~Zhang, XU~Chuchen, ZHU Jinling, CHEN Meijuan, and FANG Yiru.
\newblock Analysis of misdiagnosis of bipolar disorder in an outpatient
  setting.
\newblock \emph{Shanghai archives of psychiatry}, 30\penalty0 (2):\penalty0 93,
  2018.

\bibitem[Ke et~al.(2017)Ke, Meng, Finley, Wang, Chen, Ma, Ye, and
  Liu]{ke2017lightgbm}
Guolin Ke, Qi~Meng, Thomas Finley, Taifeng Wang, Wei Chen, Weidong Ma, Qiwei
  Ye, and Tie-Yan Liu.
\newblock Lightgbm: A highly efficient gradient boosting decision tree.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  3146--3154, 2017.

\bibitem[Khaleghi et~al.(2015)Khaleghi, Sheikhani, Mohammadi, Nasrabadi, Vand,
  Zarafshan, and Moeini]{khaleghi2015eeg}
Ali Khaleghi, Ali Sheikhani, Mohammad~Reza Mohammadi, Ali~Moti Nasrabadi,
  Safa~Rafiei Vand, Hadi Zarafshan, and Mahdi Moeini.
\newblock Eeg classification of adolescents with type i and type ii of bipolar
  disorder.
\newblock \emph{Australasian physical \& engineering sciences in medicine},
  38\penalty0 (4):\penalty0 551--559, 2015.

\bibitem[Khodayari-Rostamabad et~al.(2010)Khodayari-Rostamabad, Reilly, Hasey,
  MacCrimmon, et~al.]{khodayari2010diagnosis}
Ahmad Khodayari-Rostamabad, James~P Reilly, Gary Hasey, Duncan MacCrimmon,
  et~al.
\newblock Diagnosis of psychiatric disorders using eeg data and employing a
  statistical decision model.
\newblock In \emph{2010 Annual International Conference of the IEEE Engineering
  in Medicine and Biology}, pages 4006--4009. IEEE, 2010.

\bibitem[Knott et~al.(2001)Knott, Mahoney, Kennedy, and Evans]{knott2001eeg}
Verner Knott, Colleen Mahoney, Sidney Kennedy, and Kenneth Evans.
\newblock Eeg power, frequency, asymmetry and coherence in male depression.
\newblock \emph{Psychiatry Research: Neuroimaging}, 106\penalty0 (2):\penalty0
  123--140, 2001.

\bibitem[Liao et~al.(2017)Liao, Wu, Huang, Cheng, and Liu]{liao2017major}
Shih-Cheng Liao, Chien-Te Wu, Hao-Chuan Huang, Wei-Teng Cheng, and Yi-Hung Liu.
\newblock Major depression detection from eeg signals using kernel
  eigen-filter-bank common spatial patterns.
\newblock \emph{Sensors}, 17\penalty0 (6):\penalty0 1385, 2017.

\bibitem[Mumtaz et~al.(2017)Mumtaz, Xia, Ali, Yasin, Hussain, and
  Malik]{mumtaz2017electroencephalogram}
Wajid Mumtaz, Likun Xia, Syed Saad~Azhar Ali, Mohd Azhar~Mohd Yasin, Muhammad
  Hussain, and Aamir~Saeed Malik.
\newblock Electroencephalogram (eeg)-based computer-aided technique to diagnose
  major depressive disorder (mdd).
\newblock \emph{Biomedical Signal Processing and Control}, 31:\penalty0
  108--115, 2017.

\bibitem[Mumtaz et~al.(2018)Mumtaz, Ali, Yasin, and Malik]{mumtaz2018machine}
Wajid Mumtaz, Syed Saad~Azhar Ali, Mohd Azhar~Mohd Yasin, and Aamir~Saeed
  Malik.
\newblock A machine learning framework involving eeg-based functional
  connectivity to diagnose major depressive disorder (mdd).
\newblock \emph{Medical \& biological engineering \& computing}, 56\penalty0
  (2):\penalty0 233--246, 2018.

\bibitem[Puthankattil and Joseph(2012)]{puthankattil2012classification}
Subha~D Puthankattil and Paul~K Joseph.
\newblock Classification of eeg signals in normal and depression conditions by
  ann using rwe and signal entropy.
\newblock \emph{Journal of Mechanics in Medicine and Biology}, 12\penalty0
  (04):\penalty0 1240019, 2012.

\bibitem[Strik et~al.(1995)Strik, Dierks, Becker, and Lehmann]{strik1995larger}
WK~Strik, T~Dierks, T~Becker, and D~Lehmann.
\newblock Larger topographical variance and decreased duration of brain
  electric microstates in depression.
\newblock \emph{Journal of Neural Transmission/General Section JNT},
  99\penalty0 (1-3):\penalty0 213--222, 1995.

\bibitem[Tas et~al.(2015)Tas, Cebi, Tan, H{\i}zl{\i}-Sayar, Tarhan, and
  Brown]{tas2015eeg}
Cumhur Tas, Merve Cebi, Oguz Tan, Gokben H{\i}zl{\i}-Sayar, Nevzat Tarhan, and
  Elliot~C Brown.
\newblock Eeg power, cordance and coherence differences between unipolar and
  bipolar depression.
\newblock \emph{Journal of affective disorders}, 172:\penalty0 184--190, 2015.

\bibitem[WHO(2016)]{who2016}
WHO.
\newblock Global health estimates 2016: Disease burden by cause, age, sex, by
  country and by region, 2000-2016. geneva, world health organization.
\newblock 2016.

\bibitem[Yannick et~al.(2019)Yannick, Hubert, Isabela, Alexandre, Jocelyn,
  et~al.]{yannick2019deep}
Roy Yannick, Banville Hubert, Albuquerque Isabela, Gramfort Alexandre, Faubert
  Jocelyn, et~al.
\newblock Deep learning-based electroencephalography analysis: a systematic
  review.
\newblock \emph{arXiv preprint arXiv:1901.05498}, 2019.

\end{thebibliography}

\documentclass{midl} 

\usepackage{mwe} 
\jmlrproceedings{MIDL}{Medical Imaging with Deep Learning}
\jmlrpages{}
\jmlryear{2019}
\jmlrworkshop{MIDL 2019 -- Extended Abstract Track}

\title[ML with EEG for precise diagnosis of depression]{Machine learning with electroencephalography features for precise diagnosis of depression subtypes}

\midlauthor{\Name{Marie Zelenina\nametag{$^{1}$}} \Email{marie.zelenina@gmail.com}
\AND
\Name{Diana Maria Pinto Prata\midlotherjointauthor\nametag{$^{1,2,3}$}} \Email{diana.prata@kcl.ac.uk}\\
\addr $^{1}$ Instituto de Biofisica e Engenharia Biomedica, Faculdade de Ciencias da Universidade de Lisboa, Portugal\\
\addr $^{2}$ Institute of Psychiatry, Psychology and Neuroscience, King's College London, UK\\
\addr $^{3}$ Centre for Psychological Research and Social Intervention (CIS), University Institute of Lisbon (ISCTE-IUL), Portugal
}
\begin{document}

\maketitle

\begin{abstract}
Depression is a common psychiatric disorder, which causes significant patient distress. Bipolar disorder is characterized by mood fluctuations between depression and mania. Unipolar and bipolar depression can be easily confused because of similar symptom profiles, but their adequate treatment plans are different. Therefore, a precise data-driven diagnosis is essential for successful treatment. In order to aid diagnosis, research applied machine learning to brain imaging data, in particular to electroencephalography (EEG), with accuracies reaching 99.5\% (unipolar vs. healthy) or 85\% (bipolar vs. healthy). However, these results arise from small training sets, without validation on independent data, and thus have a high risk of inflated accuracies due to data over-fitting. We propose to use a bigger corpus of realistic clinical data for training and testing and improve classification with microstates features, which can assess the function of large-scale brain networks.
\end{abstract}

\begin{keywords}
Neuroscience, machine learning, EEG, depression, bipolar.
\end{keywords}

\section{Introduction}
Depressive disorders are the 4th leading cause of disability worldwide \cite{who2016}. Successful treatment of depression depends on a precise diagnosis, so it is important to develop classification tools that allow not only to distinguish depressed patients from healthy controls, but also to identify subtypes of depression. Especially urgent is the differentiation between unipolar and bipolar depression because they require different treatment plans, and misdiagnosis can delay recovery. It is, however, a demanding task, because unipolar and bipolar depression share a similar symptom profile \cite{hui2018analysis}. This makes a data-driven computer-aided precise diagnosis of depression subtypes based on biomarkers, rather than symptoms, an attractive approach.
Most attempts to aid in diagnosis of depression with machine learning tools applied to neuroimaging data have been centred around Magnetic Resonance Imaging (MRI) \cite{gao2018machine}. MRI benefits from advantages in spatial composition of the data, but has disadvantages in terms of low temporal composition, high cost, and requirement of more extensively trained research personnel, compared to EEG. Finally, fMRI measures brain activity indirectly, using blood oxygenation measures as a proxy of brain activity, while EEG measures the electrical activity of the brain directly. This makes EEG an attractive brain imaging tool for neuropsychiatric disorders.

\subsection{Computer-aided diagnosis of unipolar depression with EEG}

In the past 20 years, the classification of unipolar depression based on EEG features attracted significant research effort \cite{acharya2018automated,acharya2015novel,ahmadlou2012fractality,ahmadlou2013spatiotemporal,bairy2017automated,cukic2018eeg,faust2014depression,hosseinifard2013classifying,knott2001eeg,liao2017major,mumtaz2017electroencephalogram,mumtaz2018machine,puthankattil2012classification}. Notably, some of these attempts reach an accuracy of up to 99.5\%  \cite{faust2014depression}. However, all those results were achieved on small datasets (22 to 90 recordings) obtained in non-naturalistic highly controlled research settings and were not replicated in independent clinical samples, which raises doubts in scalability of results and their translatability to realistic clinical settings. 

\subsection{Computer-aided diagnosis of bipolar vs. unipolar depression with EEG}
So far, there were few attempts to differentiate unipolar from bipolar depression with machine learning tools applied to EEG data \cite{khodayari2010diagnosis,erguzel2015wrapper,erguzel2016artificial}. A maximum likelihood approach based on the combination of factor analysis models achieved an average correct diagnosis rate (prediction accuracy) of around 85\% \cite{khodayari2010diagnosis}. A support vector machine classifier achieved a prediction accuracy of 80.19\% \cite{erguzel2015wrapper}. Finally, an artificial neural network achieved a prediction accuracy of 83.87\% \cite{erguzel2016artificial}. All those attempts suffered from the same problems with scalability \& translatability as unipolar classification research. 

Given the therapeutic importance of precise diagnosis of bipolar vs. unipolar depression, further efforts are required to improve the classification precision. We suggest contributing to those efforts by adding microstates features, which can tap into the function of large-scale brain networks, and have shown great potential as biomarkers in neuropsychiatry research.

\section{Proposed methods}
\subsection{Data}
We intend to use a large EEG sample, including over 650 unipolar and over 250 bipolar depression patients and healthy controls, from multiple hospitals hospitals, for training and testing of our classifier. 90\% of this data will be used for training and testing of our model. To validate our results, we will set aside a smaller
sample (10\%) of the data, which will consist of recordings coming from a different hospital setting than the data used for training and testing.

\subsection{Features}
Inspired by previous research \cite{erguzel2015wrapper,erguzel2016artificial,khodayari2010diagnosis,tas2015eeg,cai2018pervasive,khaleghi2015eeg,cukic2018eeg,hosseinifard2013classifying,strik1995larger}, we propose the following features:
\begin{enumerate}
	\item Band features: cordance, spectral coherence, absolute power.
	\item Time domain features: mean, variance, kurtosis, skewness, Hjorth parameters.
	\item Non-linear features: entropy, C0-complezity, fractal dimension.
	\item Microstates features: duration, occurence, contribution, transition probabilities. Research has not  used microstates to classify unipolar vs. bipolar depression, although Strik et al. suggested altered microstate patters in depression \cite{strik1995larger}.
\end{enumerate}

\subsection{Machine learning}

\subsubsection{Feature selection}
We propose to use a wrapper feature selection method developed specifically for EEG data \cite{hossain2014feature}. It implements a backward search strategy, which starts with the full feature set  and iteratively removes features using correlation criteria. Because microstates have not been previously used to differentiate unipolar and bipolar depression, we intend to research microstates features separately as an additional step, with the use of analysis of variance (ANOVA), confusion matrices and single variable classifiers.

\subsubsection{Classifiers}
A recent review \cite{yannick2019deep} highlighted advantages of deep learning in applicability to EEG data because of its capacity to learn good feature representations from raw data. They discuss that deep neural networks, in particular convolutional networks, could learn features from raw or minimally preprocessed data. As such, we propose to experiment with fully connected, convolutional, and recurrent neural networks. However, we still intend to preprocess the data and perform feature selection as described above, and compare performances of pipelines with manual preprocessing and wrapper feature selection to no or minimal preprocessing and automatic feature learning with neural networks. 

Additionally, we propose to experiment classifiers such as Logistic Regression, Linear Discriminant Analysis, K-Nearest Neighbors, Decision Trees, including the recent highly popular LightGBM \cite{ke2017lightgbm}, Naive Bayes, and Support Vector Machines. The final choice of the algorithm will be done empirically. For training and testing, we will split the data with the 80-20 ratio. To maximize the amount of data available for training, we will use leave-one-out cross-validation.

\subsubsection{Evaluation techniques}
We will evaluate our model using AUC-ROC (Area Under The Curve - Receiver Operating Characteristics) analysis. It is an aggregate measure of performance across all possible classification thresholds, which can be interpreted as the probability that the model ranks a random positive example more highly than a random negative example. It is scale-invariant and measures how well predictions are ranked, rather than their absolute values. Also, it is classification-threshold-invariant and measures the quality of the model’s predictions irrespective of the classification threshold. AUC-ROC values range from 0 (worst) to 1 (best). 
Additionally, we will use the classification accuracy score, which is the ratio of number of correct predictions to the total number of input samples. 
We will consider precision accuracy higher than human diagnostic error a successful validation procedure.

\pagebreak
\bibliography{zelenina19}

\end{document}